OpenStudy (anonymous):

PLZ HELP ASAP. a solution of salt and water contains 50 grams of water per 100 millimeters of the solution. If 1 mole weighs 16 grams, _____ moles of water would be present in 25 millimeter of the solution. (round of your answer to 2 decimal places

OpenStudy (anonymous):

50 grams per 100 mm 1 mole weighs 16 grams so 50/16 = 3.125 moles 3.125/100 = x/25 (3.126 moles to 100 mm = x moles to 25 mm) cross multiply 100x = 78.125 x = 78.125/100 x = .78125 moles

OpenStudy (anonymous):

u sure right

OpenStudy (anonymous):

oops...round it to 0.78

OpenStudy (anonymous):

I think so. That is what I would put if it were my paper.

OpenStudy (anonymous):

ok... how old r u by the way just asking

OpenStudy (anonymous):

I do not give out my age. I am old enough :)

OpenStudy (anonymous):

ok lol and thx

OpenStudy (anonymous):

anytime :)